40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Sep 15 01:14:10 2015, rana, Summary, LSC, need to check LSC Whitening switch logic ... again 
    Reply  Wed Sep 16 15:04:33 2015, ericq, Summary, LSC, DC PD Whitening Board Fixed 
Message ID: 11597     Entry time: Tue Sep 15 01:14:10 2015     Reply to this: 11606
Author: rana 
Type: Summary 
Category: LSC 
Subject: need to check LSC Whitening switch logic ... again 

Tonight we noticed that the REFL_DC signal has gone bipolar, even though the whitening gain is 0 dB and the whitening filter is requested to be OFF.

We should check out the switch operation of several ofthe LSC channels in the daytime - where is the procedure for this diagnostic posted?

ELOG V3.1.3-